OpenStudy (anonymous):

Can someone please help me? :( Solve the system of equations that includes a circle and a line algebraically. Find the solutions where the conic equations intersect each other{x^2+y^2-4y=20 3x+4y=5

OpenStudy (crabbyoldgamer):

How would you normally solve a system of equations?

OpenStudy (anonymous):

I tried both elimination and plugging in one variable. But for both methods I always end up having only one variable. I'm still lost :(

OpenStudy (anonymous):

one less variable*

OpenStudy (crabbyoldgamer):

So you found out what y was in terms of x, using the second equation?

OpenStudy (crabbyoldgamer):

If so, tell me what y = in terms of x.

OpenStudy (anonymous):

Well I saw that there was a 4y in the first equation. so for the second equation I shifted the values to get 4y=5-3x. When I plugged it in I had x^2+y^2+2x-5-3x)=20

OpenStudy (crabbyoldgamer):

You can't do that. Because there is also a y^2

OpenStudy (anonymous):

Yeah, I don't understand how I can get rid of both y^2 and y

OpenStudy (crabbyoldgamer):

That is, you also have to deal with that y^2

OpenStudy (crabbyoldgamer):

Well, if 4y = 5-3x, then what does y =?

OpenStudy (anonymous):

I need to divide 4 to both sides?

OpenStudy (crabbyoldgamer):

of course

OpenStudy (crabbyoldgamer):

See, you know how to do it. You just panicked, and thought you didn't know how.

OpenStudy (anonymous):

Oh, I see now.

OpenStudy (crabbyoldgamer):

So then what does y =?

OpenStudy (crabbyoldgamer):

See?

OpenStudy (anonymous):

y=-3/4x+5/4. Then I plug in. I see now. I see now.

OpenStudy (crabbyoldgamer):

The math might be messy, but it's easy once you see that.

OpenStudy (crabbyoldgamer):

OK

OpenStudy (crabbyoldgamer):

You're welcome.

OpenStudy (anonymous):

Yes. Thank you very much:) I appreciate the help